A friend of yours doing research brags that she ha
Answer #1

Given that, 10 .00 picograms is iron (II) phosphate ; Fe3O8P2

1 .00 pg = 1e-12 g

10 .00 picograms *1*10^-12 g/ 1.00 picograms = 1.0*10^-11 g

Number of moles in 10 .00 picograms   or 1.0*10^-11 g

= amount in g/ molar mass

= 1.0*10^-11 g Fe3O8P2/ molar mass

= 1.0*10^-11 g/ 357.48 g/ mole

= 2.80*10^-14 moles

Now convert mole into molecules

= 2.80*10^-14 moles * 6.23*10^23 molecule / 1 mole

= 1.68*10^10 molecules

1 molecule of Fe3O8P2 contains = 8 atom O

1.68*10^10 molecules will contain

= 1.68*10^10 molecules *8 atom O /1 molecule of Fe3O8P2

= 1.344*10^11 atoms O